OpenStudy (anonymous):

Find the derivative of the function using the product rule (x)(sqrt(x^2+5))

OpenStudy (anonymous):

u'v+uv'

OpenStudy (anonymous):

you'll need the chain rule for the sqrt(x^2+5) term as well

OpenStudy (anonymous):

\[\sqrt{{x ^2}+5}+x ^{2}/\sqrt{x ^{2}+5}\]

OpenStudy (anonymous):

yep that looks right

OpenStudy (anonymous):

i got that but can i simplify it anymore?

OpenStudy (anonymous):

yeah you can multiply top and bottom by sqrt(x^2+5)

OpenStudy (anonymous):

the first term I mean

OpenStudy (anonymous):

(2x^2+5)/(sqrt(x^2+5))

OpenStudy (anonymous):

i mean i have to substitute my x in for the equation of the tangent line is it a good idea to just leave that like that?

OpenStudy (anonymous):

less computation if you simplify it a little further

OpenStudy (anonymous):

\[x ^{2}/\sqrt{x ^{2}+5}\]so i added it and got

OpenStudy (anonymous):

sorry 2x^2 +5 for numerator

OpenStudy (anonymous):

sorry 2x^2 +5 for numerator